whats the slope and y-intercept of the line 3x+2y=2
Answer by mananth(16946)

3x+2y=2
write in slope intercept form
2y=-3x+2
y=(-3/2)x+1
compare with y= mx+b
m=-3/2 the slope
and y intercept = 1